The sum of two numbers is -19 and their difference is -5. Write a system of equations and solve it using a graph.
I have one equation as y=-19 + (x-5) and can't come up with the other one.

Answer by Alan3354(69443)

The sum of two numbers is -19 and their difference is -5. Write a system of equations and solve it using a graph.
I have one equation as y=-19 + (x-5) and can't come up with the other one.
---------------------
x + y = -19
x - y = -5